75600J of thermal energy is needed to raise the temperature of a 2.0kg block of ice removed from a freezer at −18∘C to its melting point. Calculate the specific heat capacity.
1
Expert's answer
2020-06-29T14:06:38-0400

The thermal energy required to heat a 2-kg block from -17°C to 0°C can be expressed as


"Q=cm(t_f-t_i),\\\space\\
c=\frac{Q}{m(t_f-t_i)}=\frac{75600}{2(0-(-18))}=2100\space\frac{\text{J}}{\text{kg°C}}"